# Revisiting In-context Learning Inference Circuit in Large Language Models

### v3

@@ B.1 Data Distribution Requirement Explained by Hidden State Multiplex
Figure 12: The 3 operating magnitudes on Pythia 6.9B with various pre-training steps. Left: Step 1, Input Text Encode; Middle: Step 2, Semantics Merge; Right: Step 3, Feature Retrieval and Copy, measured by the correct induction head numbers.
@@
During the training, (1) When the label space is small, the trained attention heads only need to extract the projected length of the query on each label’s eigen-subspace. For each label, such operation has a parameters’ analytical solution with (encoding kernel) $W_{Q}^{\top}W_{K}=I$ and (decoding transformation) $W_{O}W_{V}=o_{y}^{\top}e_{y}$, where $o_{y}$ is the label token’s output embedding, and $e_{y}$ is the label’s eigen-subspace. From such an operation, theoretically, one layer can handle at most $|\mathcal{H}|$ labels, where $|\mathcal{H}|$ is the head amounts. While, considering the sparsity of these eigen-subspaces, such an upper bound can be increased to $d^{\prime}|\mathcal{H}|$ by multiplex one head to decode an orthogonal group of labels with orthogonal eigen-subspaces of $\mathcal{E}=[e_{y_{1}};e_{y_{2}};\dots;e_{y_{d^{\prime}}}]$ and orthogonal output embedding $\mathcal{O}=[o_{y_{1}};o_{y_{2}};\dots;o_{y_{d^{\prime}}}]$, where $d^{\prime}$ is the inner dimension of attention head, with decoding transformation $W_{O}W_{V}=\mathcal{O}^{\top}\mathcal{E}$. (2) When the label space expands [^12] Notice that such a situation can also occur when the orthogonality between eigen-subspaces or output embeddings is lost. A common situation is that the variance of the cluster increases, creating confusion within the decoding space of the attention head., the decoding transformation can not distinguish all the clusters since the $W_{O}W_{V}$ is low-rank. Driven by the training loss, the model can choose to transform the encoding kernel to focus on catching the most similar label tokens with the query, and copy the label tokens’ information back to the query. As a result, one attention head can catch at most $d^{\prime}$ groups of label tokens mapped collinearly by the encoding kernel (note that this set of labels may not appear simultaneously in the context, so confusion can be avoided), and the common space of these label words become the induction subspace shown in §4.2.
@@
Appendix C Can LMs Segment ICL Prompt?
The experiments in §3.1 imply that LMs conduct effective segmentation on ICL-styled inputs, enabling LMs to block the interference of preceding demonstrations in the input text encoding operation. Here, as a prototypical discussion, we confirm the existence of segmentation and then reveal that such segmentation can be done in very early layers from an attention operation focusing on some specific segmentation tokens in the inputs.
@@ Appendix C Can LMs Segment ICL Prompt?
Furthermore, we empirically demonstrate that delimiters have significant saliency towards ICL accuracies in Table 5 (upper), experimented by removing them from prompt templates. Interestingly, the trial to completely remove these delimiters from the prompt yielded almost random results, even though these inputs still conform to the primary form of ICL. A reliable reason can be that: The missing delimiter interferes with the encoding operation (Step 1) on both demonstrations and queries, so that completely disrupts the ICL process.
The scale of such a segmentation operation can surprise one since more than half of the heads focus on the segmenting operation as shown in Fig. 14 (Right). However, as an assumption, we want to argue that dividing the input text into local segmentation is a crucial step in language modeling, so, functionally, LM has sufficient motivation to focus on segmenting during the training process. Moreover, based on the above principles, as long as the delimiter appears periodically at appropriate positions and can be captured by attention heads (only in the structured parts of the prompt template), as shown in Table 5 (lower), the delimiter can be designed to any token. While we still recommend natural delimiters without semantics in the template design.
@@ Appendix D Semantics Merge is Non-selective on Location
From the results, no significant statistical differences between these two types of tokens can be observed, suggesting that the semantics copy process, which is identified as Step 2 of our circuit, is not selective on the token types. However, even if we demonstrate that the model cannot exhibit selectivity in the copy positions within the current prompt, a potential direction for future research is to investigate whether any forerunner tokens can enhance or weaken this copy process. Given that this copying is known to be related to label denoising (§4.1), exploring this and carefully designing these forerunner tokens could significantly benefit the control of ICL behavior.
@@ Appendix G Degradation Analysis: In-weight Learning without Ground-truth Label in Context
In this section, we demonstrate that our circuit, especially the induction processing of Step 3, can still explain the inference behavior in the IWL scenario, but is not as robust as in the ICL scenario. Moreover, our experiments demonstrate that in the IWL scenario, induction heads are detrimental to the model’s predictions, which is consistent with the inference drawn from the results in the main text. Also, some inference phenomena in the IWL scenario can be explained by our conclusions. While, we also illustrate through a counterexample that under the condition of IWL, even the induction heads work in a negative way, but the demonstrations can still enhance the inference.
Experimental results in this section require us to consider new inference dynamics for IWL scenarios, although we believe the discussion for ICL scenarios in this paper is robust. We demonstrate that such separate consideration of ICL and IWL is reasonable and responsible by highlighting the essential differences between ICL and IWL settings, as shown in previous works (Chan et al., 2022; Reddy, 2024).
@@ G.1 How Accurate can Induction Heads Explain IWL Inference?
Figure 19: The JS divergence between the predicted output by the induction attention score (Eq. 19) and real output on ICL and IWL settings. The lowest 5 data of each layer is selected.
@@ G.2 Do Induction Heads Contribute to IWL Processing?
In detail, compared to the results of unfiltered inputs (Table 1), we highlight 4 major phenomena, and explain them within our framework: (1) The baseline accuracies (line 1) in IWL settings are lower than in normal settings, even worse than random prediction. This is easy to understand given the frequency bias introduced in §5.3. (2) As we gradually suppress the induction heads (line 5), the accuracies increase, which is consistent with the inference in the main text of this paper: in IWL setting, induction heads can not find and copy any correct label-related information in the context, and only copy the label information presented in the context noisily. Therefore, after removing it from the inference process, the accuracies increase and approach a 0-shot level. These results confirm our main claim of this section: induction heads are detrimental to IWL inference. (3) Interestingly, removing the copying processing (line 4) unexpectedly enhances the accuracy. We infer that the input texts of different labels under the same task still have a certain background-value similarity of the encoding at the forerunner token. Such similarities are propagated into the label tokens by Step 2, enhancing the attention flow from labels to the query’s forerunner token in Step 3, causing a decrease in accuracy. (4) Removing the demonstration text encoding (line 2) decreases the accuracy to around 00. According to our discussion about Step 1, in this case, forerunner tokens cannot receive information from the demonstration text, leading an equivalent sequence for the later layers without demonstration text like $[s_{1}][y_{1}][s_{2}][y_{2}]\dots[s_{k}][y_{k}][x_{q}][s_{q}][y_{q}]$. Given that the forerunner token $[s_{\cdot}]$ is consistent throughout the input, later layers are likely to conduct induction only on the forerunner token, whose subsequent tokens do not contain the correct label, leading to a 00 accuracy. This is also a critical rebuttal to previous work (Wang et al., 2023), which believes the label tokens directly collect the information of input text: if that is the case, then disconnecting the links from demonstration texts to the forerunner tokens will not have such a significant disturbance on the accuracy.
@@ G.4 Discussions: The Possibility of Different Inference Dynamics in IWL Scenarios
Our explanation for this counterexample is: it can be considered that in the IWL settings, some different inference dynamics contribute to the prediction. Such difference is natural, since even if the ICL and IWL input data share a consistent format, they are fundamentally distinct, and sometimes even antagonistic, as shown in previous works (Chan et al., 2022; Reddy, 2024), which find that toy Transformers are difficult to perform well on both types of data with a same set of parameters. While, large models may allow for the coexistence of multiple inference dynamics, as discussed in §5.2, making LLMs able to yield better performance on both inference scenarios above.
@@ H.1 Attention Head Statistics
We count the marked count of each attention head as Forerunner Token Head (Fig. 22232425) / Correct Induction Head (Fig. 26272829) by every data sample from each dataset on each model.
Forerunner Token Head Statistics. We plot the distributions of the marked Forerunner Token Heads towards correct and wrong labels, where: there are observable morphological differences in figures across different datasets, while the forerunner token heads marked on the correct and incorrect labels of the same dataset are almost identical. The detailed data confirms our conclusion in §4.1.

### v4

@@ B.1 Data Distribution Requirement Explained by Hidden State Multiplex
Figure 12: The 3 operating magnitudes on Pythia 6.9B with various pre-training steps. Left: Step 1, Input Text Encode. Middle: Step 2, Semantics Merge. Right: Step 3, Feature Retrieval and Copy, measured by the correct induction head numbers.
@@
During the training, (1) When the label space is small, the trained attention heads only need to extract the projected length of the query on each label’s eigen-subspace. For each label, such operation has a parameters’ analytical solution with (encoding kernel) $W_{Q}^{\top}W_{K}=\mathrm{I}$ and (decoding transformation) $W_{O}W_{V}=o_{y}^{\top}e_{y}$, where $o_{y}$ is the label token’s output embedding, and $e_{y}$ is the label’s eigen-subspace. From such an operation, theoretically, one layer can handle at most $|\mathcal{H}|$ labels, where $|\mathcal{H}|$ is the head amounts. While, considering the sparsity of these eigen-subspaces, such an upper bound can be increased to $d^{\prime}|\mathcal{H}|$ by multiplex one head to decode an orthogonal group of labels with orthogonal eigen-subspaces of $\mathcal{E}=[e_{y_{1}};e_{y_{2}};\dots;e_{y_{d^{\prime}}}]$ and orthogonal output embedding $\mathcal{O}=[o_{y_{1}};o_{y_{2}};\dots;o_{y_{d^{\prime}}}]$, where $d^{\prime}$ is the inner dimension of attention head, with decoding transformation $W_{O}W_{V}=\mathcal{O}^{\top}\mathcal{E}$. (2) When the label space expands [^12] Notice that such a situation can also occur when the orthogonality between eigen-subspaces or output embeddings is lost. A common situation is that the variance of the cluster increases, creating confusion within the decoding space of the attention head., the decoding transformation can not distinguish all the clusters since the $W_{O}W_{V}$ is low-rank. Driven by the training loss, the model can choose to transform the encoding kernel to focus on catching the most similar label tokens with the query, and copy the label tokens’ information back to the query. As a result, one attention head can catch at most $d^{\prime}$ groups of label tokens mapped collinearly by the encoding kernel (note that this set of labels may not appear simultaneously in the context, so confusion can be avoided), and the common space of these label words become the induction subspace shown in §4.2.
@@
Appendix C Can LMs Segment ICL Inputs?
@@
The experiments in §3.1 imply that LMs conduct effective segmentation on ICL-styled inputs, enabling LMs to block the interference of preceding demonstrations in the input text encoding operation. Here, as a prototypical discussion, we confirm the existence of the segmentation operation and then reveal that such segmentation can be done in very early layers from an attention operation focusing on some specific segmentation tokens in the inputs.
@@
Furthermore, we empirically demonstrate that delimiters have significant saliency towards ICL accuracies in Table 5 (upper), experimented by removing them from prompt templates. Interestingly, the trial to completely remove these delimiters from the inputs yielded almost random results, even though these inputs still conform to the primary form of ICL. A reliable reason can be that: The hiatus of the delimiter interferes with the encoding operation (Step 1) on both demonstrations and queries, which completely disrupts the ICL process.
The scale of such a segmentation operation can surprise one, since more than half of the heads focus on the segmenting operation as shown in Fig. 13 (Right). However, as an assumption, we want to argue that dividing the input text into local segmentation is a crucial step in language modeling, so, functionally, LM has sufficient “motivation” to focus on segmenting by the pre-training objective. Moreover, based on the above principles, as long as the delimiter appears periodically at appropriate positions and can be captured by attention heads (only in the structured parts of the prompt template), as shown in Table 5 (lower), the delimiter can be designed to any token. While we still recommend natural delimiters without semantics in the template design.
@@ Appendix D Semantics Merge is Non-selective on Location
From the results, no significant statistical differences between these two types of tokens can be observed, suggesting that the semantics copy process, which is identified as Step 2 of our circuit, is not selective on the token types. However, even if we demonstrate that the model cannot exhibit selectivity in the copy positions within the current input, a potential direction for future research is to investigate whether any forerunner tokens in the inputs can enhance or weaken this copy process. Given that this copying is known to be related to label denoising (§4.1), exploring and carefully designing these forerunner tokens could significantly benefit the control of ICL behavior.
@@ Appendix G Degradation Analysis: In-weight Learning without Ground-truth Label in Context
In this section, we demonstrate that our circuit, especially the induction processing of Step 3, can still explain the inference behavior in the IWL scenario, but is not as robust as in the ICL scenario. Moreover, our experiments demonstrate that in the IWL scenario, induction heads are detrimental to the model’s predictions, which is consistent with the conclusion drawn from the results in the main text. Also, some inference phenomena in the IWL scenario can be explained by our conclusions. While, we also illustrate through a counterexample that under the condition of IWL, even if the induction heads work negatively, the demonstrations can still enhance the inference.
Experimental results in this section require us to consider new inference dynamics for IWL scenarios, although we believe the discussion for ICL scenarios in this paper is robust. We demonstrate that such separate consideration of ICL and IWL is reasonable and responsible by highlighting the essential differences between ICL and IWL settings, as also shown in previous works (Chan et al., 2022; Reddy, 2024).
@@ G.1 How Accurate can Induction Heads Explain IWL Inference?
Figure 19: The JS divergence between the predicted output by the induction attention score (Eq. 18) and real output on ICL and IWL settings. Since not all attention heads are induction heads, the lowest 5 data of each layer is selected.
@@ G.2 Do Induction Heads Contribute to IWL Processing?
In detail, compared to the results of unfiltered inputs (Table 1), we highlight 4 major phenomena, and explain them within our framework: (1) The baseline accuracies (line 1) in IWL settings are lower than in normal settings, even worse than random prediction. This is easy to understand given the frequency bias introduced in §5.3. (2) As we gradually suppress the induction heads (line 5), the accuracies increase, which is consistent with the conclusion in the main text of this paper: In IWL setting, induction heads can not find and copy any correct label-related information in the context, and only copy the label information presented in the context noisily. Therefore, after removing it from the inference process, the accuracies increase and approach a 0-shot level. These results confirm our main claim of this section: Induction heads are detrimental to IWL inference. (3) Interestingly, removing the copying processing (line 4) unexpectedly enhances the accuracy. We infer that the input texts of different labels under the same task still have a certain background-value similarity of the encoding at the forerunner token. Such similarities are propagated into the label tokens by Step 2, enhancing the attention flow from labels to the query’s forerunner token in Step 3, causing a decrease in accuracy. (4) Removing the demonstration text encoding (line 2) decreases the accuracy to around 00. According to our discussion about Step 1, in this case, forerunner tokens cannot receive information from the demonstration text, leading an equivalent sequence for the later layers without demonstration text like $[s_{1}][y_{1}][s_{2}][y_{2}]\dots[s_{k}][y_{k}][x_{q}][s_{q}][y_{q}]$. Given that the forerunner token $[s_{\cdot}]$ is consistent throughout the input, later layers are likely to conduct induction only on the forerunner token, whose subsequent tokens do not contain the correct label, leading to a 00 accuracy. This is also a critical rebuttal to previous work (Wang et al., 2023), which believes the label tokens directly collect the information of input text: If that is the case, then disconnecting the links from demonstration texts to the forerunner tokens will not have such a significant disturbance on the accuracy.
@@ G.4 Discussions: The Possibility of Different Inference Dynamics in IWL Scenarios
Our explanation for this counterexample is: It can be considered that in the IWL settings, some different inference dynamics contribute to the prediction. Such difference is natural, since even if the ICL and IWL input data share a consistent format, they are fundamentally distinct, and sometimes even antagonistic, as shown in previous works (Chan et al., 2022; Reddy, 2024), which find that toy Transformers are difficult to perform well on both types of data with a same set of parameters. While, large models may allow for the coexistence of multiple inference dynamics, as discussed in §5.2, making LLMs able to yield better performance on both ICL and IWL inference scenarios.
@@ H.1 Attention Head Statistics
We count the marked times of each attention head as Forerunner Token Head (Fig. 22232425) / Correct Induction Head (Fig. 26272829) by all data samples from each dataset on each model.
Forerunner Token Head Statistics. We plot the distributions of the marked Forerunner Token Heads towards correct and wrong labels: There are observable morphological differences in figures across different datasets, while the forerunner token heads marked on the correct and incorrect labels of the same dataset are almost identical. The detailed data confirms our conclusion in §4.1.